OpenStudy (anonymous):

Find the product, (z1*z2) and the quotient (z1/z2). Express your answers in polar form. (Express θ in radians.) z1 = cos π/4+ i sin π/4 z2 = cos 3π/4+ i sin 3π/4

OpenStudy (anonymous):

z1 = cis pi/4 z2 = cis 3pi/4 z1*z2 = r1r2 cis angle1+angle2 = 1*1 cis 3+1 pi/4 = 1 cis pi z1/z2 = r1/r2 cis angle1-angle2 = 1/1 cis 1-3 pi/4 = 1 cis -pi/2

OpenStudy (anonymous):

thanks for the quick reply. But what is "cis"?

OpenStudy (anonymous):

cis is short-hand for cos + i sin. r cis theta = r cos theta + i r sin theta

OpenStudy (anonymous):

oh OK. thank you for the help!